### v1

# MIRAGE: Assessing Hallucination in Multimodal Reasoning Chains of MLLM

Abstract
Multimodal hallucination in multimodal large language models (MLLMs) restricts the correctness of MLLMs. However, multimodal hallucinations are multi-sourced and arise from diverse causes. Existing benchmarks fail to adequately distinguish between perception-induced hallucinations and reasoning-induced hallucinations. This failure constitutes a significant issue and hinders the diagnosis of multimodal reasoning failures within MLLMs. To address this, we propose the MIRAGE benchmark, which isolates reasoning hallucinations by constructing questions where input images are correctly perceived by MLLMs yet reasoning errors persist. MIRAGE introduces multi-granular evaluation metrics: accuracy, factuality, and LLMs hallucination score for hallucination quantification. Our analysis reveals that (1) the model scale, data scale, and training stages significantly affect the degree of logical, fabrication, and factual hallucinations; (2) current MLLMs show no effective improvement on spatial hallucinations caused by misinterpreted spatial relationships, indicating their limited visual reasoning capabilities; and (3) question types correlate with distinct hallucination patterns, highlighting targeted challenges and potential mitigation strategies. To address these challenges, we propose Logos, a method that combines curriculum reinforcement fine-tuning to encourage models to generate logic-consistent reasoning chains by stepwise reducing learning difficulty, and collaborative hint inference to reduce reasoning complexity. Logos establishes a baseline on MIRAGE, and reduces the logical hallucinations in original base models. MIRAGE will be publicly available.
1 Introduction
Multimodal large language models (MLLMs) gpt-4o ; gemini ; grok3 achieve advancements in multimodal perception llava ; internvl ; qwen2vl , as evidenced by standard benchmarks mathvision ; mme ; chartqa ; scienceqa ; okvqa . Recent studies further enhance their reasoning capacities through post-training qvq ; virgo ; o1 ; cot ; vic . However, two critical challenges remain, i.e., erroneous visual perception that fabricates non-existent content, and defective logical reasoning yielding inconsistent conclusions. These multi-source hallucinations (stemming from distinct perceptual and cognitive origins) fundamentally limit the practical utility.
To quantitatively measure hallucination in MLLMs, several multimodal benchmarks have been applied to detect and measure multimodal hallucination in object recognition mme ; pope ; mmvp ; hallusionbench ; seedbench or academic reasoning mathvista aspects. Existing benchmarks mme ; pope ; mmvp ; hallusionbench ; seedbench ; mathvista attempt to measure hallucinations via object recognition or academic tasks. However, two critical gaps remain. First, current evaluations fail to distinguish between different types of hallucinations, i.e., perception-induced hallucinations caused by inaccurate visual understanding and reasoning-induced hallucinations stemming from logical flaws, making it difficult to pinpoint errors. Second, most benchmarks focus on validating the content of answers or intermediate steps, while lacking fine-grained evaluation of the reasoning process in terms of perception and logic, thereby hindering the ability to trace error propagation patterns. This absence of hierarchical analysis spanning answer-level outputs, step-level intermediate results, and thought-level reasoning logic prevents systematic diagnosis of reasoning failures. Addressing these gaps is essential for building trustworthy MLLMs.
Figure 1: Data distribution and structure of MIRAGE benchmark. (a) shows the classes and amounts of questions. (b) shows the data structure of each question in MIRAGE. And (c) shows the multimodal reasoning hallucination types we explored.
To address these challenges, we propose MIRAGE, a diagnostic benchmark specifically designed to isolate reasoning-induced hallucinations in MLLMs. As shown in Fig. 1, MIRAGE contains 1,329 questions where MLLMs demonstrate accurate visual perception but exhibit defective reasoning. Each question provides three-tier annotations: final answers, intermediate reasoning steps and claims, and ground-truth reasoning chains, enabling precise tracking of hallucination propagation in multimodal reasoning processes. To comprehensively assess reasoning hallucination in MLLMs, MIRAGE proposes three evaluation metrics, i.e., accuracy assessment measuring overall answer correctness, factuality assessment verifying the correctness in intermediate steps and claims, and LLMs Hallucination Score to assess hallucination from the whole reasoning chain level. By evaluating MLLMs on MIRAGE from different levels, we aim to answer three critical research questions. First, how reasoning hallucinations compromise MLLM robustness and correlate with answer accuracy. Second, whether specific question types induce distinct hallucination patterns unique to multimodal reasoning. Finally, the efficacy of current mitigation methods against reasoning-specific hallucinations.
We conduct extensive experiments on MIRAGE, leading to several key insights. First, the model scale, data scale, and training stages of MLLMs significantly influence the severity of logical, fabrication, and factual hallucinations. Second, these factors offer limited improvements in addressing spatial hallucinations, which are primarily caused by misinterpretations of spatial relationships—highlighting the limited visual reasoning capabilities of current MLLMs and their inability to benefit from straightforward scaling. Third, we observe strong correlations between question types and specific patterns of reasoning hallucination, underscoring critical challenges and suggesting targeted mitigation strategies. These findings offer valuable guidance for the future development of more reliable and reasoning-aware MLLMs.
Building on the insight that increasing the probability of logic-consistent reasoning chains reduces specific logical hallucinations, we propose Logos, which integrates curriculum reinforcement fine-tuning (CRFT) for training and collaborative hint inference for testing. During training, CRFT with online reward filtration (ORF) gradually increases question difficulty while dynamically selecting high-reward samples, guiding the model toward accurate and logic-consistent reasoning. At the testing stage, collaborative hint inference provides topic- and question-specific hints from LLMs, reducing reasoning complexity for optimized models. Experiments demonstrate that Logos significantly reduces reasoning hallucinations and achieves strong performance on both MIRAGE and standard benchmarks mathvista . In conclusion, the contribution of this paper can be summarized as follows:
  • We propose MIRAGE, the first benchmark for evaluating multimodal reasoning hallucinations in MLLMs. It isolates reasoning hallucinations with tasks where inputs are correctly perceived but reasoning errors persist, and introduces multi-level metrics for comprehensive assessment: accuracy, factuality, and LLMs hallucination score.
  • Our findings reveal that the model scale, data scale, and training stages of MLLMs, and highlight critical challenges and mitigation for specific hallucination types. These findings will provide insights for future MLLM development.
  • We propose Logos, a baseline method of MIRAGE to encourage model for logic-consistent reasoning via curriculum reinforcement fine-tuning and collaborative hint inference. Logos reduces the multimodal logical hallucination and improves the answer accuracy.
Table 1: Comparison of MIRAGE with existing benchmarks. “MCQ” means multiple-choice questions, “A” means answers, “D” means multimodal input descriptions, ‘R” means full reasoning chains, and “S” means intermediate results. means multiple reasoning chains. MIRAGE offers superior annotation coverage and assessment capabilities in reasoning hallucination assessment.
Benchmarks Dataset Properties Hallucination Assessment Usage
Taxonomy Scale Annotation Intermediate Steps Chains
POPE pope Object 18K A - Object Hallu
MMVP mmvp Object 300 A - MCQ
HallusionBench hallusionbench Object 1,129 A - Illusion
MME mme Various 2,374 A - General VQA
SEEDBench seedbench Obj+Act 19K A - MCQ
MathVista mathvista Math 1,000 A - Math Reasoning
OmniBench omnibench Various 1,142 A/D - MCQ
MME-CoT mme-cot Various 1,130 A/D/S Steps General CoT
MIRAGE (Ours) Various 1,329 A/D/R/S Steps+Hints Reason Hallu
2 Related Work
Reasoning Multimodal Large Language Models. Reasoning MLLMs can be roughly divided into three groups. First is the prompt-based reasoning method cot ; mllmcot ; compositional to guide MLLMs by in-context learning incontext . Second is the plan-based method AGoT ; BDoG ; llamaberry , which uses searching methods mcts ; prm to explore optimal reasoning chains. And last is learning-based method by supervised fine-tuning (SFT) r1-onevision or reinforcement learning (RL) deepseekmath ; visual-rft ; mm-eureka . RL methods generalize better by optimizing with their high-reward predictions instead of fixed ground-truths. Hence, we build Logos on RL for hallucination mitigation, uniquely focusing on dynamic training difficulty adjustment.
Multimodal Hallucination Evaluation. Existing MLLMs still suffer from multimodal hallucination, where generated text either contradicts the visual input or deviates from correct logical reasoning. To assess the hallucination and its effect in MLLMs, recent works measure the accuracy degradation among object perception pope , illusion hallusionbench ; autohallusion , mathematic mathvista ; mathvision ; mathverse , IQ test mmiq ; puzzlevqa ; algopuzzlevqa , and general multimodal abilities seedbench ; mme . While existing benchmarks have advanced multimodal evaluation, they often conflate perception-induced hallucinations with reasoning-induced ones, making it challenging to diagnose reasoning failures. In contrast, MIRAGE focuses on reasoning hallucinations by isolating reasoning errors from correctly perceived inputs, providing multi-level metrics for assessment.
Multimodal Hallucination Mitigation. Hallucination evaluations provide critical insights for future mitigation strategies, which broadly fall into training-free and training-based approaches. First is training-free methods vic ; selfcorrection ; react . They use prompting SoM-LLaVA , question decomposition vic , reasoning chain reflection selfcorrection ; ji-etal-2023-towards , and ensembling woo2024ritual to mitigate perception-induced hallucination. Second is training-based mitigation methods ursa ; r1-onevision ; virgo ; yang2025mitigating ; zhang2024reflective ; yu2024hallucidoctor . These methods mitigate hallucination by data curation yu2024hallucidoctor ; lrv-instruction , SFT sarkarmitigating ; ben2023mocha ; jiang2024hallucination ; chen2023mitigating , and RL ha-dpo ; zhou2024aligning ; yu2024rlhf . Previous hallucination mitigation methods mainly target perception errors in MLLMs, with only a few early efforts addressing multimodal reasoning hallucinations vic ; ursa . Our proposed Logos focuses on reasoning hallucinations and shows that combining CRFT with CHI effectively reduces hallucinations beyond prior methods.
3 MIRAGE Dataset
3.1 Data Construction
To evaluate reasoning hallucinations, as in Table 1, we present MIRAGE, emphasizing tasks with accurate perception but challenging reasoning. MIRAGE offers multi-level annotations and rich auxiliary data for error diagnosis. As in Fig. 2, the construction involves data collection and curation.
Data Collection. To systematically evaluate multimodal reasoning capacities across diverse cognitive dimensions, the MIRAGE is constructed through rigorous selection of seven distinct taxonomies, including geometry, algebraic, arithmetic, scientific, spatial reasoning, and statistical reasoning. Based on above taxonomies, we collect the original benchmark data from both publicly available datasets and questions from Internet. Finally, the size of original dataset is roughly 18K.
Data Curation. To ensure MIRAGE isolates reasoning hallucinations and comprehensively evaluate each topic, we apply a two-step curation process, i.e., difficulty curation and balance curation. For difficulty curation, we use three open-source MLLMs qwen25vl ; internvl ; llama3 to generate image descriptions, retaining only questions where these descriptions are consistently accurate (verified by a secondary LLM deepseek-v3 ) but lead to frequent reasoning errors, aligning with our benchmark focus. For balance curation, we sample the resulting data to ensure a balanced distribution lpt ; kangdecoupling ; li2022long across seven topics, maintaining a small imbalanced rate, and resulting in a final dataset of 1,329 questions.
Figure 2: The construction and evaluation of MIRAGE. (a) shows the construction of MIRAGE.And (b) shows multi-granular evaluation metrics: accuracy, factuality, and LLMs hallucination score.
3.2 Data Annotation and Verification
Reasoning Chain Annotation. To address the lack of ground-truth reasoning chains, we propose a cost-effective automated annotation framework that optimizes both computational efficiency and output quality. Our approach follows a two-stage refinement process. Firstly, we generate initial reasoning chains using the lightweight O3-mini o1 . And then we refine these chains with a strong LLM deepsek-r1 , guided by known ground-truths. We will discuss the annotation cost in appendix.
Collaborative Verification. Next, we conduct annotation verification to ensure the correctness. To improve both verification speed and accuracy, we introduce a human-AI collaborative verification framework. Specifically, each question is independently assessed by a human expert and an MLLM grok3 for potential hallucinations in the reasoning chain. If both assessments are accurate, the reasoning chain is retained as the ground-truth. In cases of discrepancies, the human expert either guides the MLLM to correct the reasoning chains or manually provides reasoning steps if the MLLM remains inaccurate. Finally, all newly annotated chains undergo cross-checking by other experts.
Step and Claim Extraction. Finally, given the verified reasoning chain and final answer for each question, we use a state-of-the-art LLM deepseek-v3 to extract critical intermediate steps and claims via in-context learning. Specifically, for each ground-truth reasoning chain $\hat{\mathbf{y}}$, we use hand-crafted few-shot prompts to guide the LLM in selecting $\hat{K}_{s}$ representative reasoning steps $\hat{\mathbb{S}}={\hat{\mathbf{s}}1,...,\hat{\mathbf{s}}_{K_{s}}}$ and $\hat{K}_{c}$ intermediate claims ^$\hat{\mathbb{C}}={\hat{\mathbf{c}}1,...,\hat{\mathbf{c}}_{K_{c}}}$. To ensure reliability, we limit $1\leq K_{s}\leq 10$ and $1\leq K_{c}\leq 10$, preventing over-detailed and unreliable outputs. The final intermediate steps and claims are then parsed using regular expressions for fine-grained reasoning hallucination evaluation.
Auxiliary Information Annotation. MIRAGE also uses MLLM gpt-4o to annotate image descriptions and hints. This information is verified by experts and can help researchers to diagnose hallucinations.
4 MIRAGE Benchmark Evaluation
4.1 Accuracy Assessment
The accuracy is a fundamental metric since incorrect final answers often indicate reasoning chain hallucinations vic ; receval . To accommodate different question types, MIRAGE parses the final predictions and matches parsed answers $\mathbf{A}_{\text{pred}}$ with ground-truths $\mathbf{A}_{\text{gt}}$ for both multiple-choice and deterministic free-form answers. For questions with approximate answers (e.g., statistical questions without precise annotations on charts), MIRAGE calculates the relative error between predictions and ground-truths, considering answers correct if the error falls below a threshold.
4.2 Factuality Assessment
Step and Claim Factuality Evaluation. For each predicted reasoning chain $y$, MIRAGE follows the extraction pipeline in Sec. 3.2 to extract intermediate steps $\mathbb{S}=\{\mathbf{s}_{1},...,\mathbf{s}_{K_{s}}\}$ and claims =$\mathbb{C}=\{\mathbf{c}_{1},...,\mathbf{c}_{K_{c}}\}$. With given corresponding ground-truth intermidiate steps $\hat{\mathbb{S}}$ and claims $\hat{\mathbb{C}}$, we utilize an LLM deepseek-v3 and use $\{\mathbb{S},\hat{\mathbb{S}}\}$ as input, and guide the LLM to detect whether a predicted step $\mathbf{s}_{i}$ is covered in $\hat{\mathbb{S}}$ and whether a ground-truth step $\hat{\mathbf{s}}_{i}$ is stated in $\mathbb{S}$, and then predicts the binary matching results $\mathbf{M}_{s,\text{pred}}$ and $\mathbf{M}_{s,\text{gt}}$. By this formulation, MIRAGE can efficiently match free-form steps and claims for flexible factuality evalution. Finally, we calculate the step factuality score $F_{\text{step}}$ by:
$F_{\text{step}}=\frac{2\times\text{Precision}_{s}\times\text{Recall}_{s}}{\text{Precision}_{s}+\text{Recall}_{s}},$ (1)
where $\text{Precision}=\frac{|\mathbf{M}_{s,\text{pred}}=1|}{|\mathbf{M}_{s,\text{pred}}|}$ means correctly answered steps and $\text{Recall}=\frac{|\mathbf{M}_{s,\text{gt}}=1|}{|\mathbf{M}_{s,\text{gt}}|}$ means correctly matched ground-truth steps. Similarly, the claim factuality score $F_{\text{claim}}$ is defined by:
$F_{\text{claim}}=\frac{2\times\text{Precision}_{c}\times\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}}.$ (2)
Hallucination Type Detection. Besides, to qualitatively assess which kind of reasoning hallucination does an MLLM suffer in a specific question, we follow LLM-as-a-Judge fu-etal-2024-gptscore ; gu2024survey and introduce an LLM-based hallucination detector. Specifically, rather than compare two plain reasoning chains directly, MIRAGE detects the reasoning hallucination by examining extracted intermediate steps $\{\mathbb{S},\hat{\mathbb{S}}\}$ and claims {$\{\mathbb{C},\hat{\mathbb{C}}\}$, then predict the hallucination detection results by in-context learning.
4.3 LLMs Hallucination Score (LHS) Assessment
Finally, we also assess the hallucination from the whole reasoning chains. While entropy-based uncertainty estimation methods uncertainty_nature ; vl-uncertainty ; zhou2024relying ; tomani2024uncertainty can identify unreliable reasoning or information-deficient chains, they still face two limitations. First is token-level likelihood dependencies. Existing methods rely on token-level likelihood to quantify uncertainty, but it is inaccessible in black-box MLLMs gemini ; gpt-4o ; grok3 . And second is high computational cost. Accurate uncertainty assessment typically requires sampling numerous responses per query, escalating evaluation overhead. Therefore, inspired by LLM as judges li2024salad ; fu-etal-2024-gptscore ; gu2024survey ; li2024llms , we propose LLMs Hallucination Score (LHS) to simulate uncertainty estimation via multi-LLMs and multi-reference ensemble. Specifically, we first define multi-dimension scoring rules to measure the hallucination in the whole reasoning chain rather than extracted steps, including factual accuracy, logical consistency, reasoning completeness, conceptual accuracy, and strategy appropriateness. Above dimensions can be used to simulate the uncertainty in responses and formulate the scoring template $\mathbf{h}_{\text{score}}$. Our aim is to predict LHS by $M$ (e.g., 3) LLM judgers. To improve the confidence of LHS, MIRAGE leverage an LLM deepseek-v3 to rewrite $\hat{\mathbf{r}}$ by $N-1$ variants, thus formulating $N=3$ reference chains $\{\mathbf{r}_{\text{ref}}^{1},...,\mathbf{r}_{\text{ref}}^{N}\}$ Then, for each response $\mathbf{r}_{\text{ref}}^{i}$ with corresponding ground-truth $\hat{\mathbf{r}}$, both responses are integrated into the template $\mathbf{h}_{\text{score}}$ and then generate the judgement scores $\{s_{1}^{i,j},...,s_{5}^{i,j}\}$ by each $j$-th LLM judger. Finally, the LHS $\bar{s}$ of response $r$ is:
$\bar{s}=\frac{1}{M}\sum_{j=1}^{M}\frac{1}{N}\sum_{i=1}^{N}\text{mean}(\{s_{1}^{i,j},...,s_{5}^{i,j}\}).$ (3)
By accumulating responses in MIRAGE, one can calcuate the mean and standard deviation of $\bar{s}$ for a specific reasoning MLLM. Generally, lower mean indicates higher uncertainty (i.e., hallucination rate), and lower standard deviation means higher confidence of the LHS. We further conduct consistency checks on LHS using human evaluators. We randomly sample 100 responses from Gemini-2-flash and Qwen2.5-VL-7B, comparing the human evaluation from three experts. The average difference rate is 7.5%, showing the reliability of LHS for measuring reasoning hallucinations.
5 Logos: A Baseline Method of MIRAGE
5.1 Revisit Multimodal Reinforcement Fine-Tuning
As shown in Sec. 6.1, reasoning chains with correct answers generally have lower hallucination rates (e.g., logical hallucination). This suggests that reducing hallucinations in MLLMs can be approached by increasing the generation probability of logic-consistent and correct reasoning chains, aligning inherently with Group Relative Policy Optimization (GRPO). To address this, we propose the baseline method Logos for MIRAGE, leveraging GRPO to optimize MLLM (the policy model $\pi$ with parameter $\theta$). Specifically, we leverage in-context learning cot to guide $\pi$ to generate formatted response with “<think>...</think>” and “<answer>...</answer>” blocks, where the former contains the reasoning chain and the latter includes the final answer. Rather than using a separate value model to calculate advantages of responses, GRPO directly samples $G$ different responses $\{\mathbf{r}_{1},...,\mathbf{r}_{G}\}$ with given multimodal question $\mathbf{x}$. To measure the relative advantages $\{A_{1},...,A_{G}\}$, we define the reward function $\mathcal{R}$ as format reward $\mathcal{R}_{\text{fmt}}$ and accuracy reward $\mathcal{R}_{\text{acc}}$, where the former is a binary function to judge whether the $i$-th response $\mathbf{r}_{i}$ follows response format, and the latter is a binary function to judge the correctness of final answer. Then the reward of $\mathbf{r}_{i}$ is defined by $r_{i}=\mathcal{R}_{\text{fmt}}(\mathbf{r}_{i})+\mathcal{R}_{\text{acc}}(\mathbf{r}_{i})$. And the advantage $A_{i}$ is:
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},...,r_{G}\})}{\text{std}(\{r_{1},...,r_{G}\})}.$ (4)
Finally, we optimize $\pi$ via minimizing GRPO loss $\mathcal{L}_{\text{GRPO}}$ with above advantages as follows:
$\mathcal{L}_{\text{GRPO}}=-\mathbb{E}_{\{\mathbf{r}_{i}\}_{1}^{G}\sim\pi(\mathbf{x})}$ $\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\mathbf{r}_{i}|}\sum_{t=1}^{|\mathbf{r}_{i}|}\min\left(\tilde{r}_{i,t}(\theta)A_{i},\text{clip}(\tilde{r}_{i,t}(\theta),1-\epsilon,1+\epsilon)A_{i}\right),$ (5)
where $\text{where }\tilde{r}_{i,t}=\frac{\pi(\mathbf{r}_{i,t}|\mathbf{x},\mathbf{r}_{i,1:t-1})}{\pi_{\text{old}}(\mathbf{r}_{i,t}|\mathbf{x},\mathbf{r}_{i,1:t-1})}.$
Figure 3: Overview of our baseline Logos. During training, Logos adopts curriculum reinforcement fine-tuning (CRFT) with online reward filtration (ORF) to progressively increase data difficulty and filter low-impact samples. During testing, Logos introduces Collaborative Hint Inference, leveraging LLM-guided hints to simplify the reasoning process. Logos effectively reduces logical hallucination.
We remove the KL-divergence term since reasoning models have a non-negligible distribution gap with base models lmm-r1 . We will investigate the effect in the App. D.
5.2 Curriculum Reinforcement Fine-Tuning
Note that if responses from a specific training sample are all correct or all incorrect, the advantage of each response is 0, which is harmful for GRPO optimization. To reduce the difficulty and improve the training efficiency, we propose curriculum reinforcement fine-tuning (CRFT). Specifically, before the optimization, we first leverage $\pi$ to conduct $G$-round sampling, and calculate average accuracy reward $\bar{r}_{\text{acc}}=\text{mean}(\{\mathcal{R}(\mathbf{r}_{1}),...,\mathcal{R}(\mathbf{r}_{G})\})$. During the first stage, we keep questions with $\bar{r}_{\text{acc}}>0$ to ensure that $\pi$ can sample at least one reasoning chain with correct answer and logic-consistent reasoning during training, thus making the advantages non-zero for smooth optimization. Then, during each $k$-round ($k>1$) curriculum training, we repeat $G$-round sampling and keep questions with $\bar{r}_{\text{acc}}<0.5$ to ensure $\pi$ can face more difficult questions during further CRFT. Our experimental results in Sec . 6.2 illustrate the efficiency of CRFT.
Table 2: Comparison of recent state-of-the-art MLLMs on MIRAGE. Best results are bolded.
Model Type Accuracy $\uparrow$ Factuality LHS $\uparrow$
$F_{\text{step}}$uparrow $F_{\text{claim}}$uparrow
Black-Box MLLMs
Gemini-2-Flash-Thinking gemini Reasoning 47.6 51.5 50.7 0.7517$\pm$0.0168
O1 o1 Reasoning 49.7 41.3 42.7 0.6193$\pm$0.0091
Gemini-2-Flash gemini General 44.1 47.8 47.4 0.6882$\pm$0.0496
GPT-4o gpt-4o General 35.0 39.2 40.6 0.6332$\pm$0.0111
Open-sourced $\sim$72B MLLMs
Qwen2.5-VL-72B-Instruct qwen25vl General 38.8 47.4 44.6 0.7223$\pm$0.0339
InternVL-2.5-78B internvl General 29.6 39.0 36.6 0.6377$\pm$0.0325
Qwen2-VL-72B-Instruct qwen2vl General 24.5 29.7 26.2 0.4928$\pm$0.0332
QvQ-72B-Preview qvq Reasoning 31.0 46.1 45.3 0.5717$\pm$0.0597
Virgo-72B virgo Reasoning 37.4 47.1 45.0 0.6328$\pm$0.0251
Open-sourced $\sim$7B MLLMs
Qwen2.5-VL-7B-Instruct qwen25vl General 28.8 34.7 31.7 0.5996$\pm$0.0123
Qwen2-VL-7B-Instruct qwen25vl General 19.5 21.9 18.6 0.3633$\pm$0.0106
Qwen2.5-VL-7B-Instruct+VIC vic Reasoning 26.9 22.8 25.2 0.4478$\pm$0.0177
Qwen2.5-VL-7B-Instruct+Reflection selfcorrection Reasoning 26.7 40.1 33.4 0.5826$\pm$0.0124
R1-OneVision-7B internvl Reasoning 22.9 30.7 30.2 0.5098$\pm$0.0099
Mulberry-Qwen2-VL-7B mulberry Reasoning 22.6 29.2 24.4 0.4740$\pm$0.0147
InternVL-2.5-8B internvl General 20.8 31.9 26.4 0.4838$\pm$0.0156
Llama-3.2-Vision-11B llama3 General 18.7 26.9 22.3 0.4265$\pm$0.0141
Llava-CoT-11B llavacot Reasoning 17.4 26.9 22.4 0.4267$\pm$0.0140
Logos-7B (Ours) Reasoning 37.1 43.3 38.3 0.6568$\pm$0.0179
Open-sourced $\sim$3B MLLMs
Qwen2.5-VL-3B-Instruct qwen25vl General 18.8 23.1 18.8 0.3422$\pm$0.0244
Phi-3.5-Instruct phi3 General 12.9 16.6 13.8 0.3181$\pm$0.0161
Logos-3B (Ours) Reasoning 29.4 38.9 34.5 0.5840$\pm$0.0216
Figure 4: Distribution between question types and reasoning hallucination types.
Figure 5: Pearson correlation among reasoning hallucination types.
5.3 Online Reward Filtration
While CRFT effectively controls data difficulty, it may still encounter training samples where all generated responses receive identical rewards, disrupting the optimization process. To address this without compromising training efficiency, we integrate offline data filtration limo into our approach, forming Online Reward Filtration (ORF). In each iteration, for given question $\mathbf{x}$ with $G$ sampled responses ${\mathbf{r}_{1},...,\mathbf{r}_{G}}$, Logos first computes the corresponding rewards ${r_{1},...,r_{G}}$ using the predefined reward function $\mathcal{R}$. If all responses share the same reward ($r_{1}=...=r_{G}$), the question is discarded for that iteration, ensuring only diverse, meaningful samples contribute to optimization.
5.4 Collaborative Hint Inference
To further reduce reasoning hallucinations beyond training, we introduce Collaborative Hint Inference (CHI), which leverages an auxiliary LLM $\phi$ to provide context-specific guidance.
Table 3: Probability of different types of reasoning hallucinations for each model.
Model Logical Spatial Factuality Context Fabrication
Gemini-2-flash 54.66% 29.33% 39.03% 28.48% 22.32%
Qwen2.5-VL-7B 68.94% 35.35% 51.07% 30.98% 23.91%
Gemini-2-flash-thinking 47.88% 25.44% 32.85% 22.51% 18.04%
Virgo-72B 63.97% 29.18% 40.65% 32.26% 21.32%
QvQ-72B-Preview 73.37% 37.93% 47.91% 47.52% 30.19%
Table 4: Manually fixing reasoning chains experimental results on 10% sampled questions with reasoning hallucination.
Model Fix Reasoning Accuracy
GPT-4o - 12.1
GPT-4o 68.5
Qwen2.5-VL-72B - 10.4
Qwen2.5-VL-72B 72.4
Table 5: Correlation matrix of all metrics. All correlations are significant at $p<0.001$ (∗∗∗).
Metric Accuracy $F_{\text{step}}$ $F_{\text{claim}}$ LHS
Accuracy 1.000 0.864∗∗∗ 0.918∗∗∗ 0.889∗∗∗
$F_{\text{step}}$ 0.864∗∗∗ 1.000 0.975∗∗∗ 0.915∗∗∗
$F_{\text{claim}}$ 0.918∗∗∗ 0.975∗∗∗ 1.000 0.933∗∗∗
LHS 0.889∗∗∗ 0.915∗∗∗ 0.933∗∗∗ 1.000
Given a question $\mathbf{x}$, CHI first uses a predefined question classification prompt to guide $\phi$ in predicting the question type $\mathbf{c}$. Based on this type, CHI generates two structured hints: a topic-specific hint $\mathbf{h}{\text{topic}}$, reflecting the general approach for the given type $\mathbf{c}$, and a question-specific hint $\mathbf{h}{\text{question}}$, tailored to the particular content of $\mathbf{x}$. During inference, we generate response by $\mathbf{r}=\pi([\mathbf{h}_{\text{topic}},\mathbf{x},\mathbf{h}_{\text{question}}])$. The optimized MLLM can benefit from CHI and generate more accurate chains than vanilla MLLMs.
6 Experiments
We state the full experimental setup of MIRAGE evaluation and training of Logos in App. B. And in the following, we illustrate the insightful findings on MIRAGE and the effectiveness of Logos.
6.1 Empirical Analysis
Overall results. As shown in Table 2, O1 achieves the highest accuracy at 49.7, outperforming Gemini-2-flash-thinking by 2.1. However, O1 scores lower in factuality, likely due to generating shorter, more summarized reasoning chains that reduce step coverage and recall. For most models except GPT-4o and O1, step scores exceed claim scores, suggesting that generating coarse reasoning steps is easier than detailed calculation steps. The LHS scores align with step scores, confirming the reliability of MIRAGE’s metrics. Focusing on open-source Qwen-VL models, increasing parameters from 3B to 72B raises accuracy from 18.8 to 38.8. Moreover, better pretraining in Qwen2.5-VL improves both accuracy and factuality/LHS, indicating that enhanced pretraining reduces hallucinations.
[…]
Correlation between reasoning hallucination and final accuracy. Given correlations between reasoning chains and answer accuracy, we conduct preliminary study to show the impact of hallucination mitigation. We manually corrected the reasoning chains for about 10% of commonly misanswered questions by GPT-4o and Qwen2.5-VL-72B, then prompt the models to reconsider their final answers. As shown in Table 5, this correction significantly improves the answer accuracy to around 70%, confirming that reducing reasoning hallucinations directly enhances overall model performance.
Table 6: Hallucination type rates in MIRAGE benchmark questions of Qwen-7/72B with different pertaining data. Pretraining with higher quality data leads to less logical, fabrication, and factual hallucinations.
Model Logical Factuality Spatial Context Fabrication
Qwen2.5-VL-72B 47.7% 33.7% 29.2% 21.6% 16.5%
Qwen2-VL-72B 59.3% 45.4% 32.7% 32.6% 26.5%
Qwen2.5-VL-7B 64.7% 45.7% 33.4% 29.3% 25.5%
Qwen2-VL-7B 74.0% 60.6% 35.6% 42.7% 35.4%
Table 7: Hallucination type rates in MIRAGE benchmark questions of Qwen-2.5-VL. Larger Models lead to less logical, fabrication, and factual hallucinations.
Model Logical Factuality Spatial Context Fabrication
Qwen2.5-VL-72B 47.7% 33.7% 29.2% 21.6% 16.5%
Qwen2.5-VL-7B 64.7% 45.7% 33.4% 29.3% 25.5%
Qwen2.5-VL-3B 78.9% 60.1% 36.7% 37.9% 38.1%
Relation between pretraining data and hallucination types. We also explore relations between pretraining data and hallucination types. Specifically, we keep use Qwen-VL qwen2vl ; qwen25vl with different pretraining data (i.e., Qwen2-VL and Qwen2.5-VL) and compare the hallucination rates of each hallucination type. As shown in Table 6, Qwen2.5-VL models have less logical, factual, and fabrication hallucination rates than those of Qwen2-VL models. A possible explanation is that pretraining data with higher quality provides more accurate factual knowledge and reasoning chains to models, such that models can avoid logical and factuality hallucinations during inference. Nevertheless, the spatial hallucination does not significantly reduced, which indicates that current MLLMs still show weak visual reasoning capabilities.
Relation between model size and hallucination types. We also explore relations between pretraining data and hallucination types. Specifically, we keep use Qwen2.5-VL qwen25vl with different model sizes (i.e., 3B/7B/72B) and compare the hallucination rates of each hallucination type. As shown in Table 7, Larger Qwen2.5-VL models have less logical, factual, and fabrication hallucination rates than those of smaller models. A possible explanation is that models owning more model parameters have more capabilities for accurate factual knowledge and reasoning chains to models, such that models can avoid logical and factuality hallucinations during inference. Nevertheless, the spatial hallucination does not significantly reduced, which indicates that current MLLMs still show weak visual reasoning capabilities.
Correlation between question and hallucination types. We also analyze the relationship between question types and hallucination patterns, as shown in Fig. 5. Results indicate that logical hallucinations are widespread across various question types, while certain hallucination types are more closely associated with specific question types. Specifically, logical and spatial hallucinations are particularly common in logical questions, reflecting the high demands for complex reasoning and visual transformations that current MLLMs struggle with mmiq . In contrast, statistical and scientific questions tend to exhibit more factuality hallucinations, likely due to their reliance on precise knowledge retrieval. These findings highlight the specific vulnerabilities of MLLMs in handling diverse reasoning tasks.
Table 8: Ablation study of Logos-7B, where CHI means collaborative hint inference.
Method GRPO CRFT $\mathbf{h}_{\text{topic}}$ $\mathbf{h}_{\text{question}}$ Accuracy $F_{\text{step}}$ $F_{\text{claim}}$ LHS MathVista
Qwen2.5-VL-7B 28.8 34.7 31.7 0.5996 68.2
+GRPO 33.7 41.0 35.9 0.6180 70.7
+CRFT 35.7 41.8 37.3 0.6193 71.9
+$\mathbf{h}_{\text{topic}}$ 36.2 42.6 37.6 0.6335 72.2
+$\mathbf{h}_{\text{question}}$ 36.5 42.2 37.6 0.6224 72.2
+full CHI 37.1 43.3 38.3 0.6568 72.3
Qwen2.5-VL-7B + full CHI 29.0 34.9 32.1 0.6011 68.3
Correlation among hallucination types. We further analyze correlations among hallucination types using pearson coefficients. As shown in Fig. 5, logical hallucinations strongly correlate with factuality, context, and fabrication errors, likely because flawed logic often leads to context inconsistency and factual errors. Notably, spatial hallucinations, which arise from complex visual operations, show relatively low correlation with other hallucinations, suggesting they are more independent and unique to multimodal models rather than text-based LLMs. These findings highlight the need for targeted mitigation strategies for hallucination types, particularly for challenging spatial reasoning errors.
Hallucination rate comparison across models. To quantitatively assess the impact of model design and training on reasoning hallucinations, we analyzed five representative MLLMs, as shown in Table 3. QvQ-72B-Preview exhibits the highest overall hallucination rates, especially in Logical (73.37%) and Context (47.52%) categories, significantly higher than Virgo-72B, which shares the same base model but benefits from more effective fine-tuning. In contrast, Gemini-2-flash-thinking consistently shows the lowest hallucination rates, particularly in Logical (47.88%), Spatial (25.44%), and Fabrication (18.04%) categories, indicating superior robustness.
Existing solutions are not sufficient to mitigate hallucination. Training-free methods like self-reflection selfcorrection and visual inference chain vic generally degrade both accuracy and LHS on base models without sufficient reasoning capabilities (Table 2), highlighting their limitations. Similarly, SFT-based methods can improve hallucination mitigation on larger models (e.g., Virgo-72B) but often fail to enhance smaller models, suggesting that model capacity plays a critical role in the effectiveness of external supervision. More detailed analysis can be found in App. D.
6.2 Empirical Analysis of Logos
We use Logos-7B as an example and conduct an ablation study on both MIRAGE and a standard benchmark, MathVista mathvista . More in-depth analysis can be found in App. D.
Table 9: Hallucination type rates in MIRAGE benchmark questions of Qwen2.5-VL-3B/7B and corresponding Logos-3B/7B. Our proposed method leads to less logical and fabrication hallucinations.
Model Logical Factual Spatial Context Fabrication
Qwen2.5-VL-7B 64.7% 45.7% 33.4% 29.3% 25.5%
Logos-7B 49.3% 39.7% 29.9% 23.8% 15.6%
Qwen2.5-VL-3B 78.9% 60.1% 36.7% 37.9% 38.1%
Logos-3B 57.1% 47.4% 36.7% 31.8% 24.0%
Comparison with previous methods. We compare Logos-7B with other 7B models to validate its effectiveness. Compared to the base model Qwen2.5-VL-7Bqwen25vl , Logos-7B achieves an 8.3 gain in accuracy, and outperforms the base by 8.6 on $F_{\text{step}}$ and 6.6 on $F_{\text{claim}}$, approaching the performance of the larger Virgo-72Bvirgo . These results, consistent with LHS scores, indicate that Logos effectively reduces reasoning hallucinations, improving reliability across reasoning chains. Similar gains are also observed for Logos-3B, showing the compatibility of our framework across different model scales.
Whether Logos reduces reasoning hallucination or not. Finally we investigate the hallucination mitigate effect on each hallucination type. As shown in Table 9, Logos-7B reduces logical hallucination by 15.4% and fabrication hallucination by 10%. Similar results can also be found in Logos-3B. Nevertheless, we do not find significant halucination mitigation on spatial and factuality hallucination on both Logos models. A possible reason is that reinforcement learning does not introduce new knowledge and only refines the logic of reasoning chains.
Ablation Study of Each Component of Logos. We first investigate the effect of each key component in Logos. The experimental results are shown in Table 8. After adopting reinforcement learning on the base model, the accuracy on MIRAGE and MathVista achieves 33.7 and 70.7 respectively. benefiting from RL, the step score and claim score also increase to 41.0 and 35.9. After adopting CRFT, both accuracy and $F_{\text{claim}}$ further increases to 35.7 and 37.3 respectively. By further integrating CHI, Logos-7B achieves 37.1 on MIRAGE and 72.3 on Mathvista. Note that directly adopt CHI on base model does not lead to performance improvement, which further proves the findings in Sec. 6.1.
How Logos mitigates reasoning hallucination? We evaluated the impact of CRFT by comparing the accuracy of Logos-7B on 8-round sampling across training dataset, before and after training. The accuracy increases from 24.8% to 68.3%, indicating that CRFT effectively guides the model to generate correct reasoning chains. Meanwhile, we calculate the “Logical” hallucination rate in Qwen2.5-VL-7B and Logos-7B, which reduces from 57.1 to 49.3. This result shows that CRFT benefits to encourage model learning logic-consistent reasoning chains to mitigate hallucination.
7 Conclusion
We propose MIRAGE, which isolates reasoning hallucinations by questions where inputs are correctly perceived but reasoning errors persist. For analysis of reasoning hallucination, MIRAGE proposes multi-level evaluation metrics, covering different levels of the reasoning chains. Our findings reveal that the model scale, data scale, and training stages of MLLMs: (1) significantly influence the degree of logical, fabrication, and factual hallucinations; (2) show no effective improvement on spatial hallucinations caused by misinterpretations of spatial relationships, suggesting that current MLLMs exhibit weak visual reasoning capabilities and struggle to benefit from simple scaling of training resources; and 3) correlations between question types and specific reasoning hallucination patterns, highlighting critical challenges and mitigation for specific types. These findings will provide insights for future MLLM development. To address this, we propose Logos, a method using curriculum reinforcement fine-tuning and collaborative hint inference to reduce logical hallucination for higher accuracy. Logos provides a baseline and offers insights for reducing hallucinations.
Appendix A Hallucination Type Definition
In this section, we summarize the hallucination types mentioned in Sec. 4.2, which are listed in Table 10. Specifically, the multimodal reasoning hallucinations can be categorized into five distinct types, i.e., spatial hallucination, logical hallucination, factuality hallucination, context hallucination, and fabrication hallucination. The detailed descriptions are summarized in Table 10.
Table 10: Categories of multimodal reasoning hallucination investigated in MIRAGE.
  Hallucination Type Hallucination Description
  Spatial Hallucination Errors in reasoning about spatial relationships, shapes, or complex visual operations.
Logical Hallucination Errors in logical consistency or reasoning, even when surface-level facts are correct.
Factuality Hallucination Factually incorrect claims about scientific principles or established knowledge in input data.
Context Hallucination Inconsistencies between intermediate reasoning steps and final predictions.
Fabrication Hallucination Entirely invented values, entities, or relationships not in input data or real world.
Appendix B Experimental Details of MIRAGE and Logos
B.1 Experimental Setup
Implementation Details. During MIRAGE evaluation, we leverage GPT-4o [23] to judge the accuracy of final answers. And for factuality and LLMs hallucination score metrics, to reduce the cost while keeping comparable evaluation accuracy, we utilize DeepSeek-V3 [36] for both metrics, and utilize Qwen2.5-72B-Instruct [67] as well as Llama-3.1-70B-Instruct [18] for LLMs hallucination score. During training of Logos, we use Qwen2.5-VL-7B-Instruct [2] as the base model. The visual encoder is frozen to avoid catastrophic forgetting of visual perception ability [22, 8, 9]. During training, we collect 13K mathematical questions with K12-level difficulty and $\sim$1K text-only math questions from LIMO [73] as training data. The batch size is 128. For each training sample, the rollout samples $G$ is 8 by default. The initial learning rate is $1\times 10^{-6}$, both warmup strategy and cosine learning rate scheduler are adopted to stabilize training. We optimize Logos by 10 epochs using AdamW [40] during each stage. The number of CRFT stages is set to 1, and we will discuss this choice in Sec. 6.2. Benefiting from the filtration mechanism in CRFT and ORF, the total training time is less than 24 hours. All programs are constructed by PyTorch [48] toolkit and vLLM [29] framework. All the experiments are conducted on 8 NVIDIA RTX A6000 GPUs.
B.2 Baseline Models
We evaluate various reasoning and CoT-enhanced general MLLMs, including black-box MLLMs [24, 55, 23] and open-sourced MLLMs [2, 60, 5, 18, 1]. We also analyze reasoning-enhanced methods [71, 66, 56, 12, 85, 16] to explore the hallucination mitigation effectiveness.
Black-box MLLMs
include GPT-4o [23] and O1 [24] from OpenAI, as well as Gemini-2-flash and Gemini-2-flash-thinking [55] from Google. These models have shown state-of-the-art reasoning or chain-of-thought thinking capabilities in various tasks.
Open-sourced MLLMs
cover both specifically-designed reasoning MLLMs (e.g., QvQ-72B [56] and Virgo-72B [12]), and General MLLMs including Qwen2-VL [60], Qwen2.5-VL [2], InternVL-2.5 [5], Llama-3.2-Vision [18] and Phi-3.5-Instruct [1]. All these models have shown competitive inherent or chain-of-thought reasoning capabilities. Note that the parameter numbers of selected models are largely varied from 3B to 72B, ensuring that models with different scales can be analyzed in our experiments.
Reasoning-enhanced General MLLMs.
Besides, to comprehensively evaluate the capabilities of reducing reasoning hallucination, we also assess multiple training-free and training-based hallucination mitigation methods, including self-reflection [16], question decomposition [85], and supervised fine-tuning [70]. All these methods have shown effectiveness in improving reasoning capabilities.
B.3 Prompts Used in Construction and Evaluation
For the reproducibility of our work, we release the critical prompts used in MIRAGE construction and evaluation. Fig. 8 shows the prompt for extracting intermediate results (i.e., steps and claims). Fig. 9 shows the prompt of intermediate results matching results (i.e., factuality evaluation prompt). Fig. 10 shows the prompt used to detect specific reasoning hallucination types. And Fig. 11 shows the prompt used to calculate the LHS score.
B.4 Computational Resources and Time
During the first stage of C-RFT, Logos-7B uses 8 NVIDIA A6000 GPUs to train a 7B model, and the total training time is 16 hours. During the second stage (since the optimal stage number $k$ is 1, as discussed in App. D), Logos-7B uses the same 8 A6000 GPUs and requires 6 hours to complete training. During Inference, benefiting from the optimization of the vLLM [29] framework, Logos-7B only requires one A6000 GPU for inference reasoning chains. And both training and inference of Logos-3B model require fewer computational resources.
B.5 Significance Computation
To calculate the correlation and the corresponding significance value, we leverage scipy package and call pearsonr function to calculate pearson correlation coefficient with corresponding significance (i.e., p-value).
Table 11: Effect of reinforcement learning algorithm. We remove CHI and directly assess the original output of each model.
RL MIRAGE MathVista
PPO 29.9 69.3
DAPO 30.0 69.6
GRPO 35.7 71.9
Table 12: Effect of online reward filtration.
ORF MIRAGE MathVista
34.2 69.6
37.1 72.3
Table 13: Effect of the KL-divergence in Logos.
KL-Div MIRAGE MathVista
1e-2 31.0 67.0
1e-3 35.0 70.0
1e-4 36.7 71.1
0 (Logos) 37.1 72.3
Table 14: Comparison between CRFT and vanilla RL with longer training epochs.
Method Total Epochs MIRAGE MathVista
Vanilla RL 20 35.5 71.4
CRFT 10+10 37.1 72.3
Table 15: Effect of curriculum learning stage $k$.
$k$ MIRAGE MathVista
0 35.0 70.7
1 37.1 72.3
2 37.2 72.3
3 37.2 72.3
Appendix C Detailed Quantitative Results
In addition to reporting average accuracy and overall LHS score for each model, we also report accuracy in each question topic and LHS score in each dimension. The detailed per-topic accuracy comparison results are shown in Table 16, and the detailed per-dimension accuracy comparison results are shown in Table 17. Generally, all models suffer from unpromising accuracy on logical and spatial questions, which indicate that existing models still do not obtain sufficient visual pattern and relation reasoning abilities. Meanwhile, mathematic reasoning and statistical question reasoning are usually perform well on state-of-the-art MLLMs. As for the LHS score, the score of logical consistency and reasoning completeness of some unpromising models are relatively low, which indicates that previous models still struggle in insufficient reasoning capabilities and result in reasoning hallucination. These results reveal the the vulnerability in reasoning MLLMs.
Appendix D More Analysis
D.1 Qualitative Results of Manually Fixing Examples
As stated in Sec. 6.2, manually fixing hallucinations in the reasoning chains enhances overall model performance. In addition to the quantitative results, we also illustrate qualitative results from GPT-4o [23], which is shown in Fig. 6. The corrected reasoning chain (right) can guide MLLMs to predict correct answers, and the original reasoning chain with hallucination (left) still results in wrong answers.
D.2 Pearson Correlation of Hallucination Type Among Single Model
In addition to stating the overall pearson correlation coefficient in Fig. 5, we also illustrate corresponding correlation from six most representative models, i.e., O1 [24], Gemini-2-flash-thinking [55], Gemini-2-flash [55], Virgo-72B [12], QvQ-72B-Preview [56], and Qwen2.5-VL-7B [2]. As shown in Fig 7, all models show a similar correlation pattern, which is consistent with Fig. 5. These results indicate the shared vulnerability in reasoning MLLMs.
Figure 6: Qualitative results of manually fixing the reasoning hallucination in the reasoning chain and inference the refined answers. The corrected reasoning chain (right) can guide MLLMs to predict correct answers, and the original reasoning chain with hallucination (left) still results in wrong answers.
Figure 7: Pearson correlation regarding hallucination types from six most representative MLLMs. All models tend to represent a similar pattern.
D.3 Training-free Methods in MIRAGE.
We also explore some training-free methods to verify the hallucination mitigation capabilities. Therefore, we evaluate prompt-based self-reflection [16] and visual inference chain [85]. As shown in Table 2, compared to Qwen2.5-VL-7B, both methods suffer performance degradation on both accuracy and LHS. These results indicate that for models without insufficient reasoning capabilities and hallucination-defending abilities, introducing training-free methods does not help to mitigate reasoning hallucination. This observation inspires us to integrate CHI into CRFT-enhanced MLLMs rather than the base model.
D.4 Supervised fine-tuning (SFT) Methods in MIRAGE
Intuitively, reasoning hallucination could be solved by further supervised fine-tuning to integrate the correct Hence, we select several SFT-based methods on 7B-level [71, 66, 70] and 72B-level [56, 12], and evaluate on MIRAGE. As shown in Table 2, on 72B MLLMs, introducing SFT can lead to better accuracy and partly mitigate reasoning hallucinations. Nevertheless, on 7B MLLMs, only Mulberry surpasses the base model by 3.1 on accuracy, while other methods do not lead to performance improvement and hallucination mitigation. This contradiction may come from the model’s capacity for model sizes. Larger models with more inherent knowledge may be easier to mitigate hallucination via external supervision, while smaller models usually struggle with SFT.
Table 16: Accuracy comparison of each question topic in MIRAGE.
  Model Algebra Arithmetic Geometry Logical Scientific Spatial Statistical Overall
           Black-Box MLLMs
Gemini-2-Flash-Thinking [55] 56.1 66.3 53.7 33.3 41.4 26.1 55.0 47.6
O1 [24] 50.9 64.1 60.0 37.7 42.5 37.8 51.9 49.7
Gemini-2-Flash [55] 51.2 57.6 50.2 34.1 32.3 26.7 55.8 44.1
GPT-4o [23] 38.8 40.2 29.5 24.6 28.0 42.2 47.2 35.0
Open-sourced $\sim$72B MLLMs
Qwen2.5-VL-72B-Instruct [2] 44.9 50.0 37.8 29.0 24.7 32.2 49.6 38.8
InternVL-2.5-78B [5] 31.1 38.0 31.4 21.7 21.0 27.2 40.3 29.6
Qwen2-VL-72B-Instruct [60] 20.1 30.4 26.0 18.8 19.9 25.0 38.0 24.5
QvQ-72B-Preview [56] 30.1 44.6 32.1 23.9 28.0 25.0 41.1 31.0
Virgo-72B [12] 44.6 47.8 37.5 29.0 23.1 38.9 41.1 37.4
Open-sourced $\sim$7B MLLMs
Qwen2.5-VL-7B-Instruct [2] 28.0 34.8 28.6 26.1 17.2 28.9 46.5 28.8
Qwen2-VL-7B-Instruct [2] 16.3 13.0 14.6 22.5 15.6 32.8 27.1 19.5
Qwen2.5-VL-7B-Instruct+VIC [85] 30.1 31.5 26.0 23.2 17.7 24.4 39.5 26.9
Qwen2.5-VL-7B-Instruct+Reflection [16] 23.9 31.5 24.8 2.3 17.2 31.7 39.5 26.7
R1-OneVision-7B [5] 20.1 22.8 21.0 26.1 16.1 28.3 33.3 22.9
Mulberry-Qwen2-VL-7B [71] 19.7 26.1 27.0 22.5 19.9 19.4 24.8 22.6
InternVL-2.5-8B [5] 11.4 26.1 22.5 16.7 17.2 30.6 30.2 20.8
Llama-3.2-Vision-11B [18] 12.5 21.7 20.0 14.5 16.1 17.2 38.0 18.7
Llava-CoT-11B [66] 12.1 14.1 19.7 10.9 16.1 20.0 31.0 17.4
Logos-7B (Ours) 39.1 39.1 38.7 32.6 20.4 34.4 59.7 37.1
Open-sourced $\sim$3B MLLMs
Qwen2.5-VL-3B-Instruct [2] 10.7 27.2 20.6 25.4 16.1 16.1 27.1 18.8
Phi-3.5-Instruct [1] 4.1 16.3 14.0 15.2 14.5 13.9 21.7 12.9
Logos-3B (Ours) 27.3 38.0 31.4 24.6 17.7 27.2 48.0 29.4
D.5 Effect of RL algorithms
As mentioned in Sec. 5, the multiple sampling pipeline in the GRPO algorithm is naturally aligned with our hallucination mitigation proposal, i.e., encouraging models to predict along the correct chain for correct answers. To verify the effect of RL algorithms, we compare Logos-7B using GRPO [53] and that using PPO [52]. We remove all CHI stages and directly assess the effect of RL. As shown in Table 11, Logos-7B using GRPO surpasses that using PPO by 5.8 and 2.6 on MIRAGE and MathVista, respectively. These results verify our motivation and provide strong support for the Logos framework design. We also evaluate the effect of recently proposed DAPO [75], a specially-designed GRPO variant. Nevertheless, it does not lead to better performance on both benchmarks. A possible explanation is that the newly introduced constraints in DAPO lead to overfitting in the model and restrict the final performance.
Table 17: LLMs Hallucination Score (LHS) comparison of each dimension in MIRAGE.
  Model Factual Logical Reasoning Conceptual Appropriateness Overall
           Black-Box MLLMs
Gemini-2-Flash-Thinking [55] 0.7182 0.7558 0.7689 0.7349 0.7372 0.7517
O1 [24] 0.6054 0.6427 0.5793 0.6384 0.6306 0.6193
Gemini-2-Flash [55] 0.6640 0.7053 0.7265 0.6862 0.7007 0.6882
GPT-4o [23] 0.5811 0.6466 0.6777 0.6198 0.6404 0.6332
Open-sourced $\sim$72B MLLMs
Qwen2.5-VL-72B-Instruct [2] 0.6330 0.7464 0.7912 0.6793 0.7321 0.7233
InternVL-2.5-78B [5] 0.5830 0.6441 0.6700 0.6088 0.6313 0.6377
Qwen2-VL-72B-Instruct [60] 0.4665 0.5115 0.5339 0.4746 0.4774 0.4928
QvQ-72B-Preview [56] 0.5024 0.5495 0.5698 0.5368 0.5168 0.5717
Virgo-72B [12] 0.6094 0.6185 0.6437 0.6252 0.6187 0.6328
Open-sourced $\sim$7B MLLMs
Qwen2.5-VL-7B-Instruct [2] 0.5333 0.6201 0.6765 0.5786 0.6130 0.5996
Qwen2-VL-7B-Instruct [2] 0.3512 0.3960 0.4120 0.3573 0.3519 0.3633
Qwen2.5-VL-7B-Instruct+VIC [85] 0.4600 0.4746 0.4336 0.4449 0.4261 0.4478
Qwen2.5-VL-7B-Instruct+Reflection [16] 0.5658 0.6242 0.6008 0.5806 0.6117 0.5826
R1-OneVision-7B [5] 0.4565 0.5227 0.5809 0.4822 0.5070 0.5098
Mulberry-Qwen2-VL-7B [71] 0.4545 0.4819 0.5070 0.4605 0.4660 0.4740
InternVL-2.5-8B [5] 0.4515 0.4967 0.5317 0.4636 0.4757 0.4838
Llama-3.2-Vision-11B [18] 0.4014 0.4473 0.4741 0.4030 0.4066 0.4265
Llava-CoT-11B [66] 0.4050 0.4417 0.4735 0.4116 0.4267 0.4267
Logos-7B (Ours) 0.5841 0.6533 0.7052 0.6233 0.6566 0.6568
Open-sourced $\sim$3B MLLMs
Qwen2.5-VL-3B-Instruct [2] 0.3282 0.3593 0.3712 0.3279 0.3242 0.3422
Phi-3.5-Instruct [1] 0.2983 0.3443 0.3459 0.3049 0.2968 0.3181
Logos-3B (Ours) 0.5486 0.5947 0.6411 0.5600 0.5757 0.5840
D.6 Effect of KL Divergence
Since we remove the KL-divergence term in Logos training, to analyze the effect, we conduct an ablation study on the KL-divergence weight. As shown in Table 13, when gradually increasing the weight of the KL-divergence term, the accuracy on both datasets is gradually decreased. When the KL-divergence term is relatively large (e.g., 1e-2), the accuracy on MathVista is even slightly lower than the base model (68.2). A possible explanation is that the distribution of reasoning MLLM has a non-negligible gap with the corresponding base models. To mitigate original reasoning hallucination and bring inherent reasoning capabilities, one should disable the KL-divergence term to tolerate the distribution gap between two models.
D.7 Effect of online reward filtration.
Next, we explore the effect of online reward filtration and report experimental results in Table 13. After integrating ORF into training, Logos-7B surpasses the counterpart by 2.9 on MIRAGE and 2.7 on MathVista. This improvement proves the effectiveness of ORF and points the future direction for more effective RL algorithms.
D.8 Effect of curriculum learning.
We also conduct experiments to verify the necessity of the CRFT stage. Specifically, we conduct vanilla RL training on all training data and ensure the number of training epochs is equal to the total CRFT. As shown in Table 15, even using longer training, one-stage RL still falls behind Logos-7B with CRFT by 1.6 on MIRAGE accuracy and 0.9 on MathVista accuracy. These results indicate that, benefiting from multi-stage difficulty filtration, the learning efficiency of Logos is highly improved. Meanwhile, we explore the effect of the curriculum learning stage $k$ in Logos. As shown in Table 15, Logos-7B can easily obtain 37.1 on MIRAGE and 72.3 on MathVista. Further increasing $k$ only introduces a marginal improvement. These results prove the effectiveness and efficiency of CRFT. Therefore, we select $k=1$ to optimize Logos.
D.9 The Quality of Automatic Annotation
We also assess the accuracy of automatically annotated reasoning chains in different phase, which is shown in Table 18. The accuracy of O3-mini initialized reasoning chains achieves 43.9. By incorporating DeepSeek-R1 and guided by answers, the accuracy of refined reasoning chains achieves 73.7. The relative high accuracy of refined reasoning chains ensures that one can reduce the human labor to correct the reasoning chains with reasoning hallucinations.
Annotation cost.
Finally we also concern the detailed annotation cost. By using the annotation method proposed in Sec. 3.2, the total cost is $\sim$22\$. And the total human working hour is 36 hours*person. We also estimate the annotation cost using O1, which is nearly 200\$. And if all the questions are annotated by human experts, the total working hour is nearly 200 hours*person. These results show the efficiency of our annotation method.
Table 18: The reasoning annotation accuracy in each phase.
Annotation Phase Reasoning Chain Accuracy
O3-mini (init) 43.9
+DeepSeek-R1 (refine w/ answer) 73.7
Appendix E Dataset Examples
To clearly show the structure of MIRAGE, we provide detailed examples of MIRAGE. Fig. 12 shows an example of geometry questions. Fig. 13 shows an example of algebraic questions. Fig. 14 shows an example of arithmetic questions. Fig. 15 shows an example of scientific questions. Fig. 16 shows an example of spatial questions. Fig. 17 shows an example of logical questions. Fig. 18 shows an example of statistical questions.
Summary of Topic-specific Hint.
We also release the topic-specific hints used in MIRAGE and Logos. As shown in Fig. 19. The topic-specific hints include key concepts and basic rules regarding the question topics. Meanwhile, the classical reasoning process of corresponding question topics is also included in the hints.
Appendix F More Qualitative Results
We also illustrate a couple of raw outputs from some representative models, i.e., Qwen2.5-VL-7B-Instruct, Gemini-2-flash-thinking, and our Logos-7B. As shown in the Fig. 20 and 21. We find that in this example, Qwen2.5-VL-7B has consistent logic in the reasoning chain, but suffers from factual hallucination (only two 90 degree angles should be calculated). In contrast, Gemini-2-flash-thinking and Logos-7B correctly solve the question.
Appendix G Limitation
The limitation of this paper is two-fold. First, MIRAGE does not include multiple images or video question-answering problems. Hence, the hallucination from the temporal dimension and the hallucination regarding cross-image relations are not fully explored. And second, the theoretical analysis of why MLLMs suffer from reasoning hallucination is still insufficient. These limitations motivate us to conduct more in-depth exploration in the future.
Appendix H Broader Impact
The broader impact of this paper lies in advancing the reliability and accuracy of multimodal large language models (MLLMs) by systematically isolating and evaluating reasoning hallucinations. MIRAGE offers a targeted benchmark for diagnosing and mitigating reasoning errors, which is essential for applications in fields like autonomous systems, medical imaging, and scientific discovery, where accurate multimodal reasoning is critical. By revealing key weaknesses in current MLLMs, such as their struggles with complex spatial reasoning, our work encourages the development of more robust, transparent, and context-aware AI systems, ultimately promoting safer and more trustworthy AI deployment.
Figure 8: The evaluation prompt used to extract intermediate results.
Figure 9: The evaluation prompt used for factuality assessment (e.g., $F_{\text{step}}$).
Figure 10: The evaluation prompt used to detect hallucination types in reasoning chains.
Figure 11: The evaluation prompt used for LLMs hallucination score extraction.
Figure 12: The example of geometry question in MIRAGE.
Figure 13: The example of algebraic question in MIRAGE.
Figure 14: The example of arithmetic question in MIRAGE.
Figure 15: The example of scientific question in MIRAGE.
Figure 16: The example of spatial question in MIRAGE.
Figure 17: The example of logical question in MIRAGE.
Figure 18: The example of statistical question in MIRAGE.
Figure 19: The topic-specific hints used in MIRAGE.
Figure 20: Response example from Qwen2.5-VL-7B-Instruct. Red font means reasoning hallucination and corresponding judgement.
Figure 21: Response examples from Logos-7B and Gemini-2-flash-thinking.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have claimed the paper’s contributions and scope in both the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discussed the limitation in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have provided all the details (including prompts) in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will release all the data and code under CC BY-NC 4.0 license.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have stated the experimental details in both main body of the paper and the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have noted the statistical significance and state the calculation method in the appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We list the detailed experiments compute resources in the appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read and confirm the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have stated the broader impacts in the appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: A few samples in the proposed dataset come from existing benchmarks (e.g., MathVista). We have cited related works and we have listed the license of related works in the appendix.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: Our paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLMs are used as a data preprocessing tool and an evaluation tool. And we also use LLM to revise the grammar of the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.